Solve the differential equation by substitution suggested by equation. Show complete solution.



(5x+3e^y)dx + 2xe^y dy =0

1
Expert's answer
2021-11-07T17:56:00-0500

We do substitution "u=e^y,du=e^ydy" and have

(5x+3u)dx+2xdu=0;

"2x\frac{du}{dx}+5x+3u=0;\\
u'(x)+\frac{5}{2}+\frac{3}{2}\frac{u}{x}=0;
t=\frac{u}{x}- new \space variable;\\
u'(x)=t'(x)x+t;\\
t'+\frac{5}{2}+\frac{5}{2}\cdot t=0;\\"

"\frac{2}{5}\frac{dt}{t+1}=-dx;"

"\frac{2}{5}\int \frac{dt}{t+1}=-\int dx=-x+C;\\
ln|t+1|=C-\frac{5}{2}x;\\
t+1=\pm e^C\cdot e^{-\frac{5}{2}x}=C\cdot e^{-\frac{5}{2}x},where \space C:=\pm e^C"

"t=-1+C\cdot e^{-\frac{5}{2}x};\\
\frac{u}{x}=-1+C\cdot e^{-\frac{5}{2}x};\\
u=-x+C\cdot x\cdot e^{-\frac{5}{2}x};\\" "C\in R"

"y=ln(u)=ln \left(x(C\cdot e^{-\frac{5}{2}x}-1) \right)" - general solution